# Sub-cycle Nanotip Field Emission of Electrons Driven by Air Plasma Generated THz Pulses

### v1

@@ Abstract
I Introduction
The development of novel sources of high brightness femtosecond electron pulses have led to significant advances in ultrafast electron microscopy and diffraction, enabling time-resolved atomic-scale imaging of dynamic processes in materials science, chemistry, and biology. Filippetto et al. (2022); Arbouet, Caruso, and Houdellier (2018) The emission of electrons from metallic surfaces and structures is often achieved by single or multi-photon photoemission and sub-cycle field emission using near-infrared femtosecond laser pulses.Ropers et al. (2007); Schenk, Krüger, and Hommelhoff (2010); Bormann et al. (2010); Echternkamp et al. (2016) Early demonstrations of nanotip electron emission employed low pulse energy, high-repetition-rate near-infrared (NIR) pulses. Hommelhoff, Kealhofer, and Kasevich (2006); Hommelhoff et al. (2006) Recently, single-cycle pulses of THz frequency light with peak fields exceeding 100 kV/cm have been used for cold-field emission (CFE) directly from the Fermi surface of metal nanotips. Li and Jones (2016); Matte et al. (2021) The metal nanotips induce a lightning rod effect, resulting in a broadband field enhancement at the tip apex of the order where is the free space wavelength and is the tip radius. Houard et al. (2020); Behr and Raschke (2008); Podenok et al. (2006)
@@ II Methods
Electron emission experiments were conducted using a tungsten nanotip housed in a spherical cube vacuum chamber, maintained at Torr. The chamber featured optical access ports and electrical feed-throughs, with the nanotip mounted on a 3-axis translation stage for precise positioning. Matte et al. (2021) Tungsten nanotips were reliably produced with tip apex raddii between 10-20 nm using electrochemical etching.Lucier (2004) A scanning electron microscope (SEM) image of nanotips used here is shown in Fig. 1 (b). Significant reshaping of the tips was observed during previous experiments,Matte et al. (2021) leading to an estimated radius during emission between 15-35 nm. This reshaping was suggested by observed emission instability immediately after installation of a fresh tip, with the electron emission being reduced by a factor of 2 over 24 hours. At pressures of Torr, this emission instability is likely due to the presence of N2 and the formation of tungsten nitride layer on the nanotip. High fields are known to cause field-assisted etching, leading to tip reshaping and even sharpening.Rezeq, Pitters, and Wolkow (2006)
@@
III Results Discussion
The THz emission efficiency and spatial focusing was optimized using a a pyroelectric power meter (Gentec-EO THZ5B-VANTA) and a microbolometer camera located at the THz focus. The time-domain profile of the detected THz pulse is shown in Fig. 2(a), with the Fourier spectra in the inset showing a spectrum peaked at 1.46 THz. The FWHM of the focused beam profile are extracted from Fig. 2(b), with values of m and m along the x- and y-axes, respectively. As shown in Fig. 2(c), at a distance of 16 mm after the focus, the beam displays the expected doughnut shape observed in previous air-plasma experiments and attributed to a Bessel-Gauss mode of the pulse.Klarskov et al. (2013) Here, the observed asymmetry is attributed to astigmatism in the imaging system. Rasmussen et al. (2023)
@@ III Results Discussion
Assuming that the electric field can be written as , where f(t) represents a normalized time-domain trace of the electric field, the total energy can be expressed as:
@@
Using a time-domain trace obtained from air biased coherent detection (ABCD), the envelope of the field is obtained through a Hilbert transform and the integral is evaluated numerically, such that = 0.073 ps. Plugging this into Eq. 5, along with our measured values of µm, µm and measured pulse energy nJ, we estimate a peak electric field of kV/cm. This field, when applied to our nanotip, would be expected to yield electron bunches of approximately e-/pulse with peak energies exceeding 6 keV, motivating the experiment presented here.
Following these initial measurements, unexpected changes were observed in the relative timing of the probe and THz pulses during the optimization of the EO signal and crystal scanning. This suggests the presence of several pulses arriving staggered in time, likely due to different emission times in the plasma or different path lengths traveled by each point-emitter in the plasma filament. This timing behavior has been fully described in a recent paper.Paulino, Colmey, and Cooke (2024) Here further evidence of this effect is described and its impact on our observed electron emission. Most importantly to our electron emission experiments, this observation indicates that the field strength provided using the time-averaged Poynting vector above may be significantly overestimated, since the power meter cannot distinguish between a single large pulse or several smaller ones being focused to the optical axis at different points in time and space. As electron emission is an excellent probe of the quasi-static peak THz field, our data will serve as another method of field strength characterization with which to compare these estimates.
@@
The electron energy spectrum was obtained by applying retarding voltages to the Faraday cup. Measurements were taken with and without silicon wafers in the collimated THz beam path to modulate the incident field. As shown in Fig. 4(a), the maximum electron energies detected were 1100 eV, 700 eV, and 450 eV for 1, 2, and 3 silicon wafers, respectively, far below the 6 keV expected from a 800 kV/cm source. These measurements were obtained by iteratively querying our electrometer, with each current value representing an integrative measurement of charge over time. The current averages and their standard deviation were calculated across five iterations, resulting in the error bars shown.
@@
Giving us direct dependence of incident light field and wavelength on the electron energy, independent of . Using our electron energies and Eq. 6, the THz field strength may once again be estimated. As this is a proportionality relationship, it can be compared with existing electron energy vs field strength data to obtain an empirical version. Matte et al. performed a similar energy scan, using THz generated by OR in LiNbO3, and determined field strengths via EO sampling and as well as using the Poynting method, displaying excellent agreement. While we also attempted to calibrate the field strength of the plasma emitted THz pulses using EO sampling, the results of this calibration were always much too low to be considered realistic, again perhaps due to the spatio-temporal profile of the THz pulse mentioned above and the much higher bandwidth of the plasma THz pulses. In Fig. 4 (b), it can be seen that for the data of Matte et al., equation 6 is better approximated by replacing the factor of with a factor of . Using this new relationship, we then estimate the plasma generated THz peak field strengths are 110, 70, 45 kV/cm, for 1, 2 and 3 wafers respectively. Accounting for the 70% transmission of each silicon wafer, the maximum field produced by air-plasma is approximately 160 kV/cm, significantly lower than the previous estimate based on THz power of 800 kV/cm. Given the magnitude of this discrepancy, it is unlikely to be caused by the aforementioned spatio-temporal profile of the THz pulse, suggesting an overestimation of field strength using the Poynting method.
@@
Fig. 5 (a) shows the measured bunch charges for the THz field strengths investigated, and a comparison to the expectation of FN theory. Assuming an emission duration of 0.4 ps and an emission area of 10 nm2, a field enhancement factor of 6900 was estimated for a local field of 70 GV/m. This value is significantly higher than similar experiments which reported values between 800 and 3800. Li and Jones (2016); Matte et al. (2021); Vella et al. (2021)
@@
These unexpected findings, including the initial overestimation of field strength and higher than expected enhancement factor, can be explained by the presence of multiple pulses arriving staggered in time. To further understand this, the ballistic acceleration of electrons in the enhanced illuminating field was simulated. Assuming the near-field fall-off given by for a hyperboloid shaped tip,Li and Jones (2016); Miller et al. (1996) the dynamics of our highest energy electrons are modeled, emitted at the peak electric field, thus undergoing the maximum possible acceleration. In this purely 1D simulation, the electron’s position and velocity are iteratively computed and updated at each time step. The code calculates the electric field at the current position, based on spatial decay and the passing pulse, to determine the force acting on the electron. The simulated temporal evolution of the electron energy is shown in Fig. 5 (b). The field enhancement factor () importantly scales as .Kim et al. (2008) As our tip radius is not precisely known, a range of tip radii are given, with a range of enhancement factors then found. Our simulations show that, for tips between 15-35 nm in radii and assuming a incident field of 110 kV/cm estimated above, a field enhancement between to could have produced our maximum detected electron energies of keV. This range is closer to our expectation based on similar experiments. We note that although the estimate based on electron energy is directly related to the highest incident field on the tip and thus remains unaffected by the presence of multiple staggered pulses, the same cannot be said for our electron counts. In the previous FN-based estimation of , it was assumed that the tip is illuminated by a single pulse with a peak field of 110 kV/cm every 1 ms, given by repetition rate of our laser. However, if this assumption is incorrect—such as in the case where multiple pulses with weaker fields arrive more frequently, our calculated would be significantly reduced, bringing it in line with the simulated values and expectation.
IV Conclusion
@@
In conclusion, sub-cycle cold field electron emission from a tungsten nanotip using a two-color laser plasma was studied, achieving electron energies of 1.1 keV and bunch charges up to electrons per pulse. While THz peak field strengths up to 800 kV/cm were estimated using the Poynting flux method, observations were consistent with a much lower field of 160 kV/cm based on the relationship between incident field and electron energy. This discrepancy was attributed to spatio-temporal focusing effects arising from collimation in the long plasma filaments using standard off-axis parabolic mirrors, producing a flying THz focus.Paulino, Colmey, and Cooke (2024)
@@
We note that preliminary results using a plastic axicon as a collimating optic showed great promise, with a more than fourfold increase in peak THz field compared to the standard off-axis parabolic mirror, suggesting THz fields of over 600 kV/cm. By adopting this new collimation scheme for future experiments, the current electron emission record of e-/pulse is expected to be surpassed, promising bunch charges up to e-/pulse and peak energies exceeding 6 keV. These advancements position plasma-driven THz sources as a promising option for ultrafast electron emission and possible THz-driven ultrafast electron microscopy using high-brightness, coherent electron beams from metal nanotips.
@@
Benjamin Colmey: Conceptualization (equal); Formal analysis (equal); Investigation (lead); Validation (equal); Writing – original draft (lead); Writing – review and editing (lead). Rodrigo T. Paulino: Conceptualization (equal); Formal analysis (equal); Investigation (equal); Validation (equal); Writing – review and editing (equal). David G. Cooke: Conceptualization (equal); Formal analysis (equal); Funding acquisition (lead); Project administration (lead); Validation (lead); Supervision (lead); Writing – review and editing (equal).

### v3

@@ Abstract
The development of sources of high brightness femtosecond electron pulses have led to significant advances in ultrafast electron microscopy and diffraction, enabling time-resolved atomic-scale imaging of dynamic processes in materials science, chemistry, and biology. Filippetto et al. (2022); Arbouet, Caruso, and Houdellier (2018) The emission of electrons from metallic surfaces and structures is often achieved by single or multi-photon photoemission and sub-cycle field emission using near-infrared femtosecond laser pulses.Ropers et al. (2007); Schenk, Krüger, and Hommelhoff (2010); Bormann et al. (2010); Echternkamp et al. (2016) Early demonstrations of nanotip electron emission employed low pulse energy, high-repetition-rate near-infrared (NIR) pulses. Hommelhoff, Kealhofer, and Kasevich (2006); Hommelhoff et al. (2006) Recently, single-cycle pulses of THz frequency light with peak fields exceeding 100 kV/cm have been used for cold-field emission (CFE) directly from the Fermi surface of metal nanotips. Li and Jones (2016); Matte et al. (2021) The metal nanotips induce a lightning rod effect, resulting in a broadband field enhancement at the tip apex of the order where is the free space wavelength and is the tip radius. Houard et al. (2020); Behr and Raschke (2008); Podenok et al. (2006)
@@
Electron emission experiments were conducted using a tungsten nanotip housed in a spherical cube vacuum chamber, maintained at Torr. The chamber featured optical access ports and electrical feed-throughs, with the nanotip mounted on a 3-axis translation stage for precise positioning. Matte et al. (2021) Tungsten nanotips were reliably produced with tip apex radii between 10-20 nm using electrochemical etching.Lucier (2004) A scanning electron microscope (SEM) image of nanotips used here is shown in Fig. 1 (b). Significant reshaping of the tips was observed during previous experiments,Matte et al. (2021) leading to an estimated radius during emission between 15-35 nm. This reshaping was suggested by observed emission instability immediately after installation of a fresh tip, with the electron emission being reduced by a factor of 2 over 24 hours. At pressures of Torr, this emission instability is likely due to the presence of N2 and the formation of tungsten nitride layer on the nanotip. High fields are known to cause field-assisted etching, leading to tip reshaping.Rezeq, Pitters, and Wolkow (2006)
@@
The THz emission efficiency and spatial focusing was optimized using a a pyroelectric power meter (Gentec-EO THZ5B-VANTA) and a microbolometer camera located at the THz focus. The time-domain profile of the detected THz pulse is shown in Fig. 2(a), with the Fourier spectra in the inset showing a spectrum peaked at 1.46 THz. The FWHM of the focused beam profile, extracted from Fig. 2(b), are m and m along the x- and y-axes, respectively, where the uncertainties were estimated using frequency-dependent beam waist variation and camera absorption profiles. As shown in Fig. 2(c), at a distance of 16 mm after the focus, the beam displays the expected doughnut shape observed in previous air-plasma experiments and attributed to a Bessel-Gauss mode of the pulse.Klarskov et al. (2013) Here, the observed asymmetry is attributed to astigmatism in the imaging system. Rasmussen et al. (2023)
@@
Assuming that the electric field can be written as , where represents a normalized time-domain trace of the electric field, the total energy can be expressed as:
@@
Using a time-domain trace obtained from air biased coherent detection (ABCD), the envelope of the field is obtained through a Hilbert transform, yielding . ABCD is preferred for this measurement as it avoids material-dependent limitations of electro-optic crystals, such as frequency-dependent phase mismatch, dispersion, and multiple reflections within the crystal, which can lead to temporal smearing and distortions in EOS Jepsen2011. The uncertainty in the time integral of the pulse envelope, , was determined by comparing the measured temporal pulse duration to that of a transform-limited Gaussian pulse, which gives a bandwidth-limited duration of . We measured the pulse energy as using a calibrated power meter, which introduces a uncertainty, corresponding to . Using these values, the peak electric field was calculated to be , with lower bound and upper bounds of and , accounting for uncertainty. These fields, when applied to our nanotip, would be expected to yield electron bunches of approximately e-/pulse with peak energies exceeding 6 keV, motivating the experiment presented here.
Following these initial measurements, unexpected changes were observed in the relative timing of the probe and THz pulses during the optimization of the EO signal and crystal scanning. This suggests the presence of several pulses arriving staggered in time, likely due to different emission times in the plasma or different path lengths traveled by each point-emitter in the plasma filament. This timing behavior has been fully described in a recent paper.Paulino, Colmey, and Cooke (2024) Here, further evidence of this effect is described and its impact on our observed electron emission. Most importantly to our electron emission experiments, this observation indicates that the field strength provided using the time-averaged Poynting vector above may be significantly overestimated, since the power meter cannot distinguish between a single large pulse or several smaller ones being focused to the optical axis at different points in time and space. As electron emission is an excellent probe of the quasi-static peak THz field, our data will serve as another method of field strength characterization with which to compare these estimates.
@@
The electron energy spectrum was obtained by applying retarding voltages to the Faraday cup. Measurements were taken with and without silicon wafers in the collimated THz beam path to modulate the incident field. As shown in Fig. 4(a), the maximum electron energies detected were 1100 eV, 700 eV, and 450 eV for 1, 2, and 3 silicon wafers, respectively, far below the 6 keV expected from a 600 kV/cm source. These measurements were obtained by iteratively querying our electrometer, with each current value representing an integrative measurement of charge over time. The current averages and their standard deviation were calculated across five iterations, resulting in the error bars shown.
@@
Giving us direct dependence of incident light field and wavelength on the electron energy, independent of . Using our electron energies and Eq. 6, the THz field strength may once again be estimated. As this is a proportionality relationship, it can be compared with existing electron energy vs field strength data to obtain an empirical version. Matte et al. performed a similar energy scan, using THz generated by OR in LiNbO3, and determined field strengths via EO sampling and as well as using the Poynting method, displaying excellent agreement. While we also attempted to calibrate the field strength of the plasma emitted THz pulses using EO sampling, the results of this calibration were always much too low to be considered realistic, again perhaps due to the spatio-temporal profile of the THz pulse mentioned above and the much higher bandwidth of the plasma THz pulses. In Fig. 4 (b), it can be seen that for the data of Matte et al., equation 6 is better approximated by replacing the factor of with a factor of . Using this relationship, we then estimate the plasma generated THz peak field strengths are 112.4, 71.5, 46.0 kV/cm, for 1, 2 and 3 wafers respectively. Accounting for the 70% transmission of each silicon wafer, the maximum field produced by air-plasma is approximately 160 kV/cm, significantly lower than the previous estimate based on THz power of 600 kV/cm. Given the magnitude of this discrepancy, it is unlikely to be caused by the aforementioned spatio-temporal profile of the THz pulse, suggesting an overestimation of field strength using the Poynting method.
@@
Fig. 5 (a) shows the measured bunch charges for the THz field strengths investigated, and a comparison to the expectation of FN theory. Assuming an emission duration of 0.4 ps and an emission area of 10 nm2, a field enhancement factor of 6900 was estimated for a local field of 70 GV/m. This value is significantly higher than similar experiments which reported values between 800 and 3800. Li and Jones (2016); Matte et al. (2021); Vella et al. (2021). Potential explanations include variations in the work function due to surface contamination, which might alter the local field and geometry of the tip; deviations in emission area due to changing tip morphology; the presence of multiple emission sites on the tip surface; and space-charge effects, which could have influenced our observed electron energy distributions by causing additional broadening. However, these effects alone are unlikely to fully account for the observed discrepancies.
@@
Another possible explanation is the presence of multiple pulses arriving staggered in time. To further understand this, the ballistic acceleration of electrons in the enhanced illuminating field was simulated. Assuming the near-field fall-off given by for a hyperboloid shaped tip,Li and Jones (2016); Miller et al. (1996) the dynamics of our highest energy electrons are modeled, emitted at the peak electric field, thus undergoing the maximum possible acceleration. In this purely 1D simulation, the electron’s position and velocity are iteratively computed and updated at each time step. The code calculates the electric field at the current position, based on spatial decay and the passing pulse, to determine the force acting on the electron. The simulated temporal evolution of the electron energy is shown in Fig. 5 (b). The field enhancement factor () importantly scales as .Kim et al. (2008) As our tip radius is not precisely known, a range of tip radii are given, with a range of enhancement factors then found. Our simulations show that, for tips between 15-35 nm in radii and assuming a incident field of 110 kV/cm estimated above, a field enhancement between to could have produced our maximum detected electron energies of keV. This range is closer to our expectation based on similar experiments. We note that although the estimate based on electron energy is directly related to the highest incident field on the tip and thus remains unaffected by the presence of multiple staggered pulses, the same cannot be said for our electron counts. In the previous FN-based estimation of , it was assumed that the tip is illuminated by a single pulse with a peak field of 110 kV/cm every 1 ms, given by repetition rate of our laser. However, if this assumption is incorrect—such as in the case where multiple pulses with weaker fields arrive more frequently, our calculated would be significantly reduced, bringing it in line with the simulated values and expectation.
@@
In conclusion, sub-cycle cold field electron emission from a tungsten nanotip using a two-color laser plasma was studied, achieving electron energies of 1.1 keV and bunch charges up to electrons per pulse. While THz peak field strengths up to 600 kV/cm were estimated using the Poynting flux method, observations were consistent with a much lower field of 160 kV/cm based on the relationship between incident field and electron energy. This discrepancy was attributed to spatio-temporal focusing effects arising from collimation in the long plasma filaments using standard off-axis parabolic mirrors, producing a flying THz focus.Paulino, Colmey, and Cooke (2024)
@@
We note that preliminary results using a plastic axicon as a collimating optic showed great promise, with a more than fourfold increase in peak THz field compared to the standard off-axis parabolic mirror, suggesting THz fields of over 600 kV/cm. By adopting this collimation scheme for future experiments, the current electron emission record of e-/pulse is expected to be surpassed, promising bunch charges up to e-/pulse and peak energies exceeding 6 keV. These advancements position plasma-driven THz sources as a promising option for ultrafast electron emission and possible THz-driven ultrafast electron microscopy using high-brightness, coherent electron beams from metal nanotips.
@@
Benjamin Colmey: Conceptualization (equal); Formal analysis (equal); Investigation (lead); Validation (equal); Writing – original draft (lead); Writing – review and editing (lead). Rodrigo T. Paulino: Conceptualization (equal); Formal analysis (equal); Investigation (equal); Validation (equal); Writing – review and editing (equal). Gaspard Beaufort: Investigation (supporting). David G. Cooke: Conceptualization (equal); Formal analysis (equal); Funding acquisition (lead); Project administration (lead); Validation (lead); Supervision (lead); Writing – review and editing (equal).